OpenStudy (anonymous):

Multiply and simplify by factoring. 4√10,000 * 4√10,000,000

OpenStudy (anonymous):

is this the fourth root?

OpenStudy (anonymous):

\[\sqrt[4]{10000}=\sqrt[4]{10^4}=10\] \[\sqrt[4]{10000000}=\sqrt[4]{10^7}=10\sqrt[4]{10^3}\] so you get \[100\sqrt[4]{10^3}=100\sqrt[4]{1000}\]